A fair 6-sided number cube is rolled once and a fair coin is flipped once. What is the probability that on odd number is rolled and a head is flipped?
Answer by stanbon(57424)

A fair 6-sided number cube is rolled once and a fair coin is flipped once. What is the probability that on odd number is rolled and a head is flipped?
----
These events are independent so you multiply their probabilities.
P(odd on the cube) = 3/6 = 1/2
P(head on the coin) = 1/2
----
P(odd and head) = (1/2)(1/2) = 1/4
=========================================
Cheers,
Stan H.